OpenStudy (anonymous):

Find the value when x = 2 and y = 3 x^-3y^-3

OpenStudy (lgbasallote):

-2^-3(3)^-3

OpenStudy (lgbasallote):

just substitute x and y into the equation

OpenStudy (anonymous):

Substitute the values for x and y. Rewrite the equation so you have positive exponents. 1. 2(2)^0(3)^(-2)=2(1)(1/(3^2))=2(1/9)=2/9 A number to the zero exponent is 1 Negative exponents x^-2=1/x^2 just another way to represent a fraction